# Benchmarking quantum computers

### v1

@@ Abstract
Quantum computers will not outperform classical computers for all computational tasks, but there’s good reason to believe they can speed up specific structured tasks. If such a task is also (subjectively) useful, it can represent quantum utility. A challenge problem is a specific instance of such a task, crafted to be (1) feasible with a quantum computer, (2) infeasible—or at least more costly—with any other computer, and (3) useful. Plausible challenge problems are sufficiently rare and distinct that it makes sense to treat each one individually.
@@
The best candidate challenge problems involve calculating quantities that can’t feasibly be computed without a quantum computer. Although problems that are just very hard for classical computers and much easier with quantum computers are also appealing, they are moving targets. Because classical algorithms can often be improved, and the power and speed of classical computers grows steadily over time, even a 1000×1000$\times$1000 × quantum advantage could be erased within years (or even days, with rapid classical algorithm development to counter a claimed quantum advantage [98, 99, 100]). So the most compelling challenge problems are specific instances of computational problems for which the quantum computer has exponential advantage in a key resource (generally time-to-solution).
@@
Two of the best studied categories of challenge problems are integer factorization and quantum chemistry. Instances of factorization are easy to describe (e.g., “factor this 2048-bit semiprime” [7]), they scale naturally in size and difficulty, and their solutions can be verified easily and rigorously. Instances of quantum chemistry can be stated precisely (e.g., “sample from the energy eigenspectrum of the FeMo cofactor to at least chemical accuracy” [8]), but their solutions are typically only heuristically verifiable, and “hardness” is challenging to quantify systematically even though a domain expert can usually assess a given instance’s difficulty.
@@
Few sharply defined challenge problems like factoring and chemistry are known. Establishing more would be very useful, but appears to be very hard. This has fostered confusion about how soon quantum utility might be achieved. The few known challenge problems suggest that quantum utility is very difficult, and will not be achieved soon. But there is a huge gray area containing important problems that might admit quantum speedups and quantum-friendly problems that might be useful. Identifying even one additional unambiguous challenge problem from among them could change the benchmarking landscape drastically.
Discrete optimization, machine learning, and linear algebra offer promising candidate challenge problems, but it’s not clear whether useful speedups even exist in these areas. If they do, the resulting challenge problems will probably resemble quantum chemistry (difficult to systematize) more than factoring (sharp and verifiable). Factoring may be unique among known challenge problems in the concision and clarity of any stated instance.
@@
To get more good challenge problems, we may need to rely on heuristic notions of verification and difficulty. If so, consensus will only emerge from detailed analyses of the best classical algorithms’ capabilities and substantive debate between domain experts. Recent developments in quantum chemistry illustrate this process. Active debate about what instances are “hard” [101], and about which quantum advantages will be both feasible and useful [102], has proceeded in parallel with research advances in simulation algorithms and careful constant-factor resource estimates [103, 8, 104] that are necessary to define and assess quantum utility.
@@
Figure 4: Assessing quantum computer performance via capability. This figure illustrates one way to compare experimentally benchmarked performance against resource estimates for challenge problems, using a multidimensional capability metric. Challenge problems and benchmark tasks are represented by the width (some measure of the number of qubits) and depth (some measure of the number of clock cycles) of a quantum circuit that performs the task. Regions indicate the circuits performable by two real-world quantum computers—Google’s Sycamore (green) as extrapolated from results in Arute et al. [2], and an ensemble of IBM Q devices (pink) benchmarked by our group [37]—and one hypothetical quantum computer (blue) (we use a success threshold of $1/e$). Points indicate constant-factor resource estimates for three candidate challenge problems analyzed in the literature [7, 8, 9]. For these problems, width is the number of logical qubits, not accounting for logical qubits used in distillation or routing, and depth is the total number of non-Clifford operations (i.e., Toffoli and/or T gates). These metrics are somewhat crude, but indicate the rough scale of resources required for these challenge problems. We emphasize the wide gulf between that “utility” scale and current state of the art capabilities—logarithmic axes were required to compress both scales into one figure. Plots like this one could enable stakeholders to track and extrapolate the growth of quantum computer capabilities over time, toward eventual achievement of quantum utility.
@@
Resource estimation addresses the question “How big, fast, and reliable would a quantum computer need to be?” in order to solve a specific challenge problem [103, 105, 106, 107, 7, 8, 108, 109, 104, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 9]. A resource estimate for a challenge problem is a description of a minimal (known) quantum computer that would suffice to solve the challenge problem efficiently. We say “minimal” rather than “minimum” because some resources can be traded for others (e.g., memory vs speed, or space vs time).
Quantum algorithms are typically analyzed in terms of their asymptotic scaling, e.g., the manner in which the requisite quantum computational resources scale with problem size, solution accuracy, etc. Resource estimation often carries the connotation that the constant prefactors for specific problem instances are also calculated. Resource estimates are quantitative, and indicate whether theoretical quantum speedups are practically feasible. For example, recent work [131, 132] provides a strong argument that good challenge problems are unlikely to involve merely quadratic quantum speedups, even under optimistic assumptions about advances in quantum hardware and architectures and relatively pessimistic assumptions about the performance of classical hardware.
Resource estimates can be specified at various levels of abstraction and detail, and different levels of detail can motivate entirely different benchmarking paradigms. The purpose of a resource estimate, at any level of detail, is to describe the key properties and performance characteristics of a plausible quantum computer that could solve a challenge problem.
A simple logical resource estimate might be “$N$ qubits, capable of executing [specific gates], with error probability $\leq\epsilon$ per gate, running for time $T$,” derived directly from a fully-compiled algorithm known to solve the problem. Logical resource estimates for most known challenge problems mandate error probabilities of $\epsilon<10^{-15}$/gate, and error rates this low are widely believed to be achievable only via fault-tolerant QEC.
More granular resource estimates account for the unavoidable existence of “hard” and “easy” logic operations in every fault-tolerant architecture (see Box 2). They count the minimum number and cost of “hard” operations that will need to be implemented using tricks like magic-state distillation and injection [16] or code switching [133]. A more detailed approach specifies a particular QEC code and method of implementing fault-tolerant computation within that code (e.g., lattice surgery [94]).
The most sophisticated physical resource estimates provide a full architectural specification—“$N^{\prime}$ physical qubits, capable of implementing [specific fault-tolerant quantum computing scheme], with error probability $\leq\epsilon^{\prime}$ per physical gate, running for time $T^{\prime}$.” Today’s best physical resource estimates use oversimplified error models, and future resource estimates may treat the physical hardware in even greater detail in order to specify engineering specs like materials quality or temperature.
@@
Resource estimates (for specific challenge problems) tell us where the goalposts of quantum utility are. Roadmaps define the distance to those goalposts. Tracking progress toward a specific computationally useful computer requires knowing what progress looks like. A roadmap describes a sequence of increasingly capable prototype devices. Some steps in this sequence may introduce key technologies that don’t immediately increase computational power. Roadmaps enable benchmarking to recognize such steps as progress and quantify their value.
@@
Each prototype can be associated with a list of measurable specifications, e.g., gate error rates, latency times, crosstalk rates, control bandwidth, etc. A roadmap is (at minimum) a sequence of such specification lists, which trace out a path to a useful device. Real devices are enormously complex, with many specifications. To track progress along a roadmap using benchmarks, a short list of key specifications needs to be identified. Those specifications should be (i) measurable or inferrable using benchmarks, and (ii) sufficient to prove that the prototype achieves its goals within the context of the roadmap.
@@
The most relevant specifications (and thus the benchmarks necessary to infer them) will change at different points along a roadmap. For example, in the early stages of a roadmap toward a fault-tolerant quantum computer with millions of qubits [7, 8, 9], gate error rates on physical qubits may be paramount and require direct benchmarking. In later stages, after multiple fault-tolerant logical qubits have been assembled, logical qubit error rates may replace them as a key specification.
Benchmarking progress towards utility
In the context provided by a challenge problem, a resource estimate, and a roadmap, benchmarks can be constructed for the purpose of measuring the specifications of a physical quantum computer (i.e., a prototype), and identifying its position along the roadmap (see Fig. 1b). The best benchmarks for this task will vary with the scale of the prototype and the details of the roadmap.
@@
One way to measure a prototype’s performance relative to a roadmap is to use benchmarks that isolate properties of specific low-level components (e.g., error rates of physical one- and two-qubit gates) and combine them using calculations or simulations to infer the values of higher-level metrics (which can then be compared to target values) such as logical qubit error rates. This approach is implicit in fault-tolerance threshold theorems that compute high-level computational properties from low-level models. However, quantum computers are complex systems, whose emergent behavior may not be reliably captured or predicted by modeling informed by component benchmarks [37, 25, 75, 24]. Holistic benchmarks (see Fig. 2) offer a complementary approach. They can probe computational power directly, provide direct evidence of progress along a roadmap, and can be used to calibrate simulations.
@@
It is unlikely that any single scalar metric will accurately assess or incentivize progress toward quantum utility. However, holistic benchmarks can report rich multi-dimensional metrics with greater descriptive power. A particular example that we have found useful and inspiring is capability benchmarks [37]. Since a quantum computer’s computational power comes from running programs, one way to probe that power is to ask “What programs can it run?” The set of programs that a computer can run, with reasonable accuracy using reasonable time and energy, is called its capability [37, 134]. Formally, this set-valued metric is infeasible to measure or even write down because there are far too many programs. But it can be sketched or approximated using a short list of key program features. If sufficiently faithful, a sketch of capability provides intuitive heuristic answers to many questions about a quantum processor’s computational ability.
@@
Capabilities can be sketched by choosing a limited set of program features, such as circuit width and circuit depth [135, 37] or the number of “hard” gates (e.g., two-qubit gates) in a circuit [6]. We define a circuit class containing all programs with the same values of the selected features. Ideally, we seek to choose features so that a given processor will be able to successfully run every circuit in a class, or none of them. Inasmuch as this holds, a benchmark of moderate complexity can probe the processor’s ability to run sample programs from each circuit class. The result is a high-dimensional metric that can be visualized as a capability region [37] (Fig. 4) in the space defined by the program features—and compared directly to resource estimates, for challenge problems, displayed on the same plot.
@@
This specific approach has clear limitations that have not yet been overcome (What kind of programs should be run? Do good features exist? How many are needed? What defines “the same” program on different architectures? Should the programs be defined at high or low levels of abstraction, and what kinds of compilation should be allowed?). Perhaps an entirely different approach will work better. But we suspect that any successful strategy for benchmarking progress to utility will need to move beyond the single-scalar-metric paradigm, and will need to address these challenges or similar ones.
@@
An open problem for any such approach is how to find or construct genuinely representative proxy programs that behave like programs that would solve a challenge problem, but that can be scaled seamlessly to fit on any given prototype. A related open question is how these programs change when adapted to different architectures, and how to construct benchmarks that evaluate both NISQ and fault-tolerant architectures on equal footing. We foresee a significant era during which advanced NISQ devices compete with, and should be fairly compared to, early fault-tolerant devices. They may pursue very different challenge problems via very different roadmaps. Creating benchmarks that can compare progress of all architectures in a shared context seems challenging, but valuable to many stakeholders.

### v2

@@ Abstract
Quantum computers will not outperform classical computers on all computational tasks, but there’s good reason to believe they can speed up specific structured tasks. If such a task is also (subjectively) useful, then executing that task successfully can constitute quantum utility. A challenge problem is a specific instance of a computational problem that is crafted to be feasible to solve with a quantum computer but infeasible—or at least more costly—with any other computer, and is also useful to solve.
@@
The most elegant challenge problems involve calculating quantities that cannot feasibly be computed without a quantum computer. Although problems that are just very hard for classical computers are also appealing, they are moving targets. This is because classical algorithms continue to improve and the power and speed of classical computers continues to grow, so even a $1000\times$ quantum advantage could be erased within years (or even days, with rapid classical algorithm development to counter a claimed quantum advantage) [121, 122, 123, 124]. So the most compelling challenge problems are specific instances of computational problems for which quantum computers have exponential advantage in a key resource (typically time-to-solution).
@@
Two of the best studied categories of challenge problems are integer factorization and quantum chemistry. Instances of factorization are easy to describe (e.g., “factor this 2048-bit semiprime” [8, 125]), they scale naturally in size and difficulty, and their solutions can be verified easily. Instances of quantum chemistry can be stated precisely (e.g., “estimate an energy eigenvalue of the FeMo cofactor to at least chemical accuracy” [9]), but their solutions are typically only heuristically verifiable, and “hardness” is challenging to quantify systematically even though a domain expert can usually assess a given instance’s classical difficulty.
@@
Few sharply defined challenge problems have been formulated, and the existing challenge problems suggest that achieving quantum utility is very difficult. But there is a huge gray area containing important problems that might admit quantum speedups and quantum-friendly problems that might be useful to solve. This has fostered confusion about how soon quantum utility might be achieved. Discrete optimization, machine learning, and linear algebra offer promising candidate challenge problems, but it’s not clear whether any useful speedups exist in these areas. However, identifying even one additional unambiguous challenge problem from among them (e.g., via the development of a new or improved quantum algorithm) could change the benchmarking landscape drastically.
@@
To identify more challenge problems, we may need to rely on heuristic notions of verification and difficulty. If so, consensus will only emerge from detailed analyses of the best classical algorithms’ capabilities and substantive debate between domain experts. Recent developments in quantum chemistry illustrate this process. Active debate about what instances are “hard” [126], and about which quantum advantages will be both feasible and useful [127], has proceeded in parallel with advances in quantum algorithms for simulation and the careful constant-factor resource estimates [128, 9, 129] that are necessary to assess quantum utility.
@@
Figure 4: Assessing quantum computer performance via capability. This figure illustrates one way to compare experimentally benchmarked performance against resource estimates for challenge problems, using a multidimensional capability metric. Challenge problems and benchmark tasks are represented by the width (some measure of the number of qubits) and depth (some measure of the number of clock cycles) of a quantum circuit that performs the task. Regions indicate the circuits performable by two real-world quantum computers—Google’s Sycamore (green) as extrapolated from results in Arute et al. [3], and an ensemble of IBM Q devices (pink) benchmarked by our group [39]—and one hypothetical quantum computer (blue) (we use a success threshold of $1/e$). Points indicate constant-factor resource estimates for three candidate challenge problems analyzed in the literature [8, 9, 10]. For these problems, width is the number of logical qubits, not accounting for logical qubits used in distillation or routing, and depth is the total number of non-Clifford operations (i.e., Toffoli and/or T gates). These metrics are somewhat crude, but indicate the rough scale of resources required for these challenge problems (but note that the required resources could decrease with new or improved quantum algorithms). We emphasize the wide gulf between that “utility” scale and even our hypothetical 100-qubit quantum computer (with beyond state-of-the-art capabilities)—logarithmic axes were required to compress both scales into one figure. Plots like this one could enable stakeholders to track and extrapolate the growth of quantum computer capabilities over time, toward eventual achievement of quantum utility.
@@
Resource estimation addresses the question “How big, fast, and reliable would a quantum computer need to be?” in order to solve a specific computational problem [128, 130, 131, 132, 8, 9, 133, 134, 129, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 10]. A resource estimate for a challenge problem is a description of a minimal (known) quantum computer that would suffice to solve that problem. We say “minimal” rather than “minimum” because some resources can be traded for others (e.g., memory vs speed, or space vs time).
Quantum algorithms are typically first analyzed in terms of their asymptotic scaling, i.e., how the required computational resources scale with problem size, solution accuracy, etc. Resource estimation means that the constant prefactors for specific problem instances are also calculated. Resource estimates are quantitative, and indicate whether theoretical quantum speedups are practically feasible. For example, recent resource estimates [156, 157] provide a strong argument that good challenge problems are unlikely to involve merely quadratic quantum speedups, even under optimistic assumptions about quantum hardware advances and pessimistic assumptions about the performance of classical hardware.
Resource estimates can be specified at various levels of abstraction and detail, and different levels of detail could motivate the use or development of entirely different benchmarks. A simple logical resource estimate for a challenge problem might be “$N$ qubits, capable of executing [specific gates], with error probability $\leq\epsilon$ per gate, running for time TT,” derived directly from the quantum circuits of an algorithm known to solve the problem. Logical resource estimates for most existing challenge problems mandate error probabilities of $\epsilon<10^{-15}$/gate, which are widely believed to be achievable only via fault-tolerant QEC.
[…]
@@
Resource estimates for challenge problems tell us where the goalposts of quantum utility are. Roadmaps specify an engineering route to get to those goalposts—they specify how a particular quantum computing technology can plausibly progress towards quantum utility. A roadmap describes a sequence of increasingly capable prototype devices. Some steps in this sequence may introduce key technologies that don’t immediately increase computational power and/or aren’t rewarded by today’s most widely-used (computational) benchmarks. Roadmaps enable a broader view of benchmarking that recognizes and quantifies all the steps along a path to utility.
@@
Each prototype in a roadmap can be associated with a list of measurable specifications, e.g., gate error rates, latency times, crosstalk rates, control bandwidth, etc. A roadmap is (at minimum) a sequence of such specification lists, which trace out a path to a useful device. Real devices are enormously complex, with many specifications. To track progress along a roadmap using benchmarks, a short list of key specifications needs to be identified. Those specifications should be measurable, and sufficient to prove that the prototype achieves its goals within the context of the roadmap.
@@
The most relevant specifications (and thus the benchmarks necessary to infer them) will likely change at different points along a roadmap. For example, in the early stages of a roadmap toward a fault-tolerant quantum computer with millions of qubits [8, 9, 10], gate error rates on physical qubits may be paramount and require direct benchmarking. In later stages, after multiple fault-tolerant logical qubits have been assembled, logical qubit error rates may replace them as a key specification. A robust understanding of what the most important specifications are, at each stage of a particular quantum computing technology’s development, will be important for the development of benchmarks that successfully track progress towards utility.
Benchmarks that measure progress towards utility
@@
In the context provided by challenge problems, resource estimates, and a roadmap, we can assess whether a particular benchmark measures progress along that roadmap, and (if necessary) design new benchmarks that do measure this progress. An as-built prototype can be compared to an idealized prototype in a specific roadmap (see Fig. 1b) by using benchmarks that directly measure each of the key performance specifications of that prototype. This enables answering questions such as “are crosstalk error rates as low as planned?”. The benchmarks needed for this task will vary with the scale of the prototype and the details of the roadmap. As quantum computing technology progresses, the most relevant specifications for a prototype quantum computer will be increasingly formulated in terms of high-level metrics (e.g., logical qubit error rates, or circuit success probabilities). One way to measure metrics like this is to do so indirectly by measuring the properties of the lowest-level components (e.g., error rates of physical one- and two-qubit gates) and combining them using calculations or simulations to infer the values of higher-level metrics (e.g., logical qubit error rates). This approach is implicit in fault-tolerance threshold theorems that compute high-level computational properties from low-level models. However, quantum computers are complex systems, whose emergent behavior may not be reliably captured or predicted by modeling informed by component benchmarks [39, 29, 91, 28], so this indirect approach to benchmarking is unlikely to be sufficient.
Benchmarks that measure performance metrics that are specific to a particular roadmap will be essential for informing the engineering of useful quantum computers. But there is also a need to quantify progress towards utility with high-level, intuitive, and technology-agnostic performance metrics. Holistic benchmarks that probe computational power directly can provide intuitive and direct evidence of progress along a roadmap. It is unlikely that any single scalar metric of computational power will accurately assess or incentivize progress toward quantum utility. However, holistic benchmarks can report rich multi-dimensional metrics with greater descriptive power. A particular example that we have found useful and inspiring is capability benchmarking [39].
@@
Since a quantum computer’s computational power comes from running quantum programs, one way to probe that power is to ask “What quantum programs can it run?” The set of quantum programs that a computer can run, with reasonable accuracy using reasonable time and energy, is called its capability [39, 159]. Formally, this set-valued metric is infeasible to measure because there are far too many quantum programs. But it can be sketched or approximated using a short list of key program features. If sufficiently faithful, a sketch of capability provides intuitive heuristic answers to many questions about a quantum processor’s computational ability.
@@
Capabilities can be sketched by choosing a limited set of program features, such as circuit width and circuit depth [54, 39] or the number of “hard” gates in a circuit [7]. We define a circuit class containing all programs with the same values of the selected features. Ideally, we seek to choose features so that a given processor will be able to successfully run every circuit in a class, or none of them. Inasmuch as this holds, a benchmark of moderate complexity can probe the processor’s ability to run sample programs from each circuit class. The result is a high-dimensional metric that can be visualized as a capability region [39] (Fig. 4) in the space defined by the program features. These capability regions can be compared directly to resource estimates, for challenge problems, displayed on the same plot, as well as computed for future prototypes on a roadmap.
@@
This specific approach has clear limitations that have not yet been overcome (What kind of programs should be run? Do good features exist? How many are needed? What defines “the same” program on different architectures? Should the programs be defined at high or low levels of abstraction? What kinds of compilation should be allowed?). But we suspect that any successful strategy for benchmarking progress to utility will need to move beyond the single-scalar-metric paradigm, and will need to address these challenges or similar ones.
@@
An open problem for any such approach is how to find or construct genuinely representative proxy programs that behave like programs that would solve a challenge problem, but that can be scaled seamlessly to fit on any given prototype. A related open question is how these programs change when adapted to different architectures, and how to construct benchmarks that evaluate both NISQ and fault-tolerant architectures on equal footing. There might be a significant era during which advanced NISQ computers compete with, and should be fairly compared to, early fault-tolerant quantum computers. They may pursue very different challenge problems via very different roadmaps. Creating benchmarks that can compare progress of all architectures in a shared context seems challenging, but valuable to many stakeholders.